OpenStudy (anonymous):

Vertex form of: x^2+3x-(1/2)+y=0 The (1/2) is one half.

OpenStudy (anonymous):

Will be willing to metal someone.

OpenStudy (freckles):

y=a(x-h)^2+k is vertex form

OpenStudy (anonymous):

I mean't the equation, not the formula or whatever that is :p

OpenStudy (freckles):

\[y=-x^2-3x+\frac{1}{2}\]

OpenStudy (freckles):

So we are going to complete the square so we can put it in the form y=a(x-h)^2+k

OpenStudy (anonymous):

did you divide by -1 to do that?

OpenStudy (freckles):

unless you are saying you don't want to put it in vertex form now...

OpenStudy (anonymous):

I do, I just wasn't taught that today.

OpenStudy (freckles):

I subtracted everything on both sides except the y

OpenStudy (anonymous):

yeah you subtracted the -y to the other side, and then divided by -1, to keep the y positive?

OpenStudy (freckles):

Well I subtracted everything on both sides EXCEPT the y

OpenStudy (anonymous):

ohhh.

OpenStudy (freckles):

But you could subtract y on both sides then multiply both sides by -1

OpenStudy (anonymous):

but you'll still get the same outcome

OpenStudy (anonymous):

alright so would i complete the square or quadratic formula?

OpenStudy (freckles):

Well you want the vertex form which is y=a(x-h)^2+k so we need to complete the square

OpenStudy (anonymous):

Can you teach me how to complete the square? I know how to do it, but I forgot how to do it with odd numbers. :'p

OpenStudy (freckles):

\[y=(-x^2-3x)+\frac{1}{2} \\ y=-(x^2+3x)+\frac{1}{2} \\ y=-(x^2+3x+?)+\frac{1}{2}+?\]

OpenStudy (anonymous):

:o, i thought completing the square was half of three times that by 2, add it to both sides?

OpenStudy (freckles):

\[x^2+bx+(\frac{b}{2})^2=(x+\frac{b}{2})^2\]

OpenStudy (anonymous):

Jaja, sorry but my teacher hasn't taught us anything like that. x'D

OpenStudy (freckles):

That is how you complete the square...

OpenStudy (anonymous):

oh i reread the post, i get it.

OpenStudy (anonymous):

i thought the steps were different with odd numbers

OpenStudy (freckles):

So your b was 3 and you said let's add inside that ( ) (3/2)^2 right?

OpenStudy (anonymous):

right

OpenStudy (freckles):

\[x^2+3x+(\frac{3}{2})^2=(x+\frac{3}{2})^2 \]

OpenStudy (anonymous):

wait i am confused, what did you do with the (-) -> -x^2-3x

OpenStudy (freckles):

\[y=(-x^2-3x)+\frac{1}{2} \\ y=-(x^2+3x)+\frac{1}{2} \\ y=-(x^2+3x+?)+\frac{1}{2}+? \\ \text{ the ? is } (3/2)^2 \\ y=-(x^2+3x+(\frac{3}{2})^2)+\frac{1}{2}+(\frac{3}{2})^2\]

OpenStudy (anonymous):

did you do this -(x^2-3x)

OpenStudy (freckles):

well no i did this: -(x^2+3x)

OpenStudy (anonymous):

yeah typo in the ( ) x'D

OpenStudy (anonymous):

3/2 = 1.5 * 2 = 3

OpenStudy (freckles):

Why are you doing 3/2*2?

OpenStudy (anonymous):

i meant 3 divided by 2, the outcome times by 2

OpenStudy (anonymous):

or square it***

OpenStudy (anonymous):

lol my bad.

OpenStudy (freckles):

\[y=(-x^2-3x)+\frac{1}{2} \\ y=-(x^2+3x)+\frac{1}{2} \\ y=-(x^2+3x+?)+\frac{1}{2}+? \\ \text{ the ? is } (3/2)^2 \\ y=-(x^2+3x+(\frac{3}{2})^2)+\frac{1}{2}+(\frac{3}{2})^2 \\ y=-(x+\frac{3}{2})^2+\frac{1}{2}+\frac{3}{2} \cdot \frac{3}{2}\]

OpenStudy (freckles):

I wrote (3/2)^2 as 3/2*3/2 so you wouldn't think (3/2)^2 means 3/2*2

OpenStudy (anonymous):

I just went blonde for the past 15 minutes. x'D

OpenStudy (anonymous):

I got y= -(x+1.5)^2 -(5/2) (5/2) = 5 over 2(fraction)

OpenStudy (anonymous):

wait nevermind, i did it wrong.

OpenStudy (freckles):

\[\text{ These are the steps } \\ y=ax^2+bx+c \\ y=ax^2+\frac{a}{a}bx+c \text{ remember a/a=1} \\ y=a(x^2+\frac{1}{a}bx)+c \text{ I factored whatever was in front of the } \\ x^2 \text{ from both of the first two terms } \\ y=a(x^2+\frac{b}{a}x)+c \\ y=a(x^2+\frac{b}{a}x+?)+c-a \cdot ? \\ \text{ the } ? \text{ is } (\frac{b}{2a})^2 \\ y=a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)+c-a \cdot (\frac{b}{2a})^2 \\ y=a(x+\frac{b}{2a})^2+c-a \cdot \frac{b^2}{4a^2} \\ y=a(x+\frac{b}{2a})^2+c-\frac{b^2}{4a} \\ y=a(x+\frac{b}{2a})^2+\frac{4ac-b^2}{4a}\] Where the vertex is : \[(\frac{-b}{2a},\frac{4ac-b^2}{4a})\]

OpenStudy (anonymous):

I know how to find the vertex, but we got to set the quadratic equation to vertex form and with that, we get the vertex.

OpenStudy (freckles):

Well I showed you just now the process for getting both the vertex equation and the vertex form.

OpenStudy (freckles):

and i mean the vertex also

OpenStudy (anonymous):

Yes thanks, but my school teaches it in a terrible way.